A customer has asked a caterer for 60lbs of nuts, 60% of which are to be cashews. The caterer has available mixtures off 70% cashews and 45% cashews. How many pounds of each mixture should be used?
Answer by checkley71(8403)

.7X+.45(60-X)=.6*60 OR .7X+27-.45X=36 OR .25X=9 OR X=36 LB. OF 70% &
60-36=24 LB. OF 45%